OpenStudy (buttermequeasy):

Solve for x: −3|2x + 6| = −12

OpenStudy (anonymous):

|2x + 6| = 4 2x + 6 = 4 2x= -2 x = -1 or 2x+6 = -4 2x = -10 x = -5 (-1,-5)